u/u-u+1/2=4
The professor gave this question at the end of class and he did not get to go over it. At first I tried to get rid of the two on the bottom but I got stuck when it came to getting rid of all the u's.

Answer by stanbon(75887)

u/u-u+1/2=4
-----------------
(u/u) - u + (1/2) = 4

1 - u = (7/2)
u = 1 -(7/2)
u = -5/2
==============
Cheers,
Stan H.